OpenStudy (ray_sihota):

write the equation of line passing through (-3, 4) having a slope of -6. I know I am suppose to put it in y=mx+b formula 4=-6(-3)+b?

OpenStudy (bradely):

y-y1=m(x-x1) y-4=-6(x-(-3)) y-4=-6x-18 6x+y+14=0 Source: http://www.mathskey.com/question2answer/ask

OpenStudy (ray_sihota):

thank you so much